Tier1 Operations Report 2018-03-21

From GridPP Wiki
Jump to: navigation, search

RAL Tier1 Operations Report for 21st March 2018

Review of Issues during the week 14th to 21st March 2018
  • This week has seen a disproportionately large number of GGUS tickets being raised, all appearing to have xrootd as the common denominator. After extensive investigations it has been found that this issue can be tracked down to a firewall rule being recently lost and consequently packets being dropped (appropriate thanks to Rajan and John for providing the appropriate proof for this PM Tuesday). At time of writing, we believe the rule has been re-instated. As this represents a major incident it has been proposed that further investigation as to the cause should be undertaken.
Current operational status and issues
  • The problems with data flows through the site firewall being restricted is still present. New firewall equipment has been delivered and is scheduled for installation in a few weeks time.
Resolved Castor Disk Server Issues
  • gdss622 (d0t1) - genTape - Drive physically moved and re-cabled. Issue found with new cables not working, has now been resolved.
Ongoing Castor Disk Server Issues
  • None
Limits on concurrent batch system jobs.
  • CMS Multicore 550
Notable Changes made since the last meeting.
Entries in GOC DB starting since the last report.
  • None
Declared in the GOC DB
Service Scheduled? Outage/At Risk Start End Duration Reason
All Castor SCHEDULED OUTAGE 27/03/2018 10:00 27/03/2018 16:00 6 hours Outage to Castor Storage System while back-end databases are patched. During the intervention both Castor disk and tape will be unavailable.
Advanced warning for other interventions
The following items are being discussed and are still to be formally scheduled and announced.
  • Replacement (upgrade) of RAL firewall scheduled for a few weeks time.

Listing by category:

  • Echo
    • Applying minor CEPH update to fix the "backfill" bug. Increase number of placement groups and add more capacity.
  • Castor:
    • Update systems to use SL7 and configured by Quattor/Aquilon. (Tape servers done)
    • Move to generic Castor headnodes.
  • Networking
    • Extend the number of services on the production network with IPv6 dual stack. (Done for Perfsonar, FTS3, all squids and the CVMFS Stratum-1 servers).
    • Replacement (upgrade) of RAL firewall.
  • Internal
    • DNS servers will be rolled out within the Tier1 network.
  • Infrastructure
    • Testing of power distribution boards in the R89 machine room is being scheduled for some time late July / Early August. The effect of this on our services is being discussed.
Open GGUS Tickets (Snapshot during morning of meeting)
Request id Affected vo Status Priority Date of creation Last update Type of problem Subject Scope
117683 none on hold less urgent 18/11/2015 09/03/2018 Information System CASTOR at RAL not publishing GLUE 2 EGI
124876 ops on hold less urgent 07/11/2016 13/11/2017 Operations [Rod Dashboard] Issue detected : hr.srce.GridFTP-Transfer-ops@gridftp.echo.stfc.ac.uk EGI
127597 cms on hold urgent 07/04/2017 20/03/2018 File Transfer Check networking and xrootd RAL-CERN performance EGI
132589 lhcb in progress very urgent 21/12/2017 14/03/2018 Local Batch System Killed pilots at RAL WLCG
133992 atlas in progress less urgent 12/03/2018 15/03/2018 File Transfer RAL-LCG2-ECHO: No such file or directory EGI
134030 other waiting for reply less urgent 14/03/2018 20/03/2018 Other request to replicate larsoft.opensciencegrid.org cvmfs repository EGI
134049 cms in progress urgent 15/03/2018 21/03/2018 CMS_AAA WAN Access Low HC xrootd success rates at T1_UK_RAL WLCG
134086 cms waiting for reply urgent 16/03/2018 20/03/2018 CMS_Central Workflows Failures reading files T1_UK_RAL WLCG
134103 atlas in progress urgent 17/03/2018 19/03/2018 File Transfer UK RAL-LCG2-ECHO: transfer errors with "the server responded with an error 500 500-Command failed" WLCG
134119 cms in progress urgent 19/03/2018 20/03/2018 CMS_Central Workflows Issues reading files through xrootd T1_UK_RAL WLCG
134136 atlas in progress less urgent 20/03/2018 20/03/2018 File Transfer RAL-LCG2: [SE][Ls][SRM_INVALID_PATH] No such file or directory EGI
GGUS Tickets Closed Last week
Request id Affected vo Status Priority Date of creation Last update Type of problem Subject Scope
133619 cms solved top priority 21/02/2018 19/03/2018 CMS_Central Workflows T1_UK_RAL Unmerged files missing WLCG
133717 cms verified less urgent 27/02/2018 19/03/2018 CMS_Data Transfers RAL FTS3 Service: Significant Drop in Transfer Efficiency WLCG
133719 atlas solved urgent 27/02/2018 14/03/2018 File Transfer Transfers to RAL-LCG2-ECHO failing WLCG
133752 atlas solved very urgent 01/03/2018 14/03/2018 File Transfer RAL FTS service appears broken WLCG
133764 snoplus.snolab.ca solved very urgent 01/03/2018 14/03/2018 Information System BDII missing SFU information EGI
133997 lhcb verified urgent 13/03/2018 14/03/2018 File Access Bad data was encountered WLCG
134037 lhcb solved very urgent 15/03/2018 20/03/2018 File Access File access problem at RAL-LCG2 WLCG
134042 dteam solved top priority 15/03/2018 15/03/2018 Other Test Ticket. WLCG
134095 lhcb solved urgent 16/03/2018 21/03/2018 File Access Bad data encountered WLCG
134147 alice solved urgent 20/03/2018 20/03/2018 Storage Systems RAL tape SE for ALICE failing since March 14 WLCG
Availability Report
2018-03-07 100 100 100 99 100 100
2018-03-08 100 100 100 98 100 100
2018-03-09 100 100 100 98 100 100
2018-03-10 100 100 100 97 100 100
2018-03-11 100 100 100 100 100 100
2018-03-12 100 100 100 100 100 100
2018-03-13 100 98 100 100 100 100
Hammercloud Test Report

Key: Atlas HC = Atlas HammerCloud (Queue RAL-LCG2_UCORE, Template 841); CMS HC = CMS HammerCloud

Day Atlas HC CMS HC Comment
2018/03/14 94 99
2018/03/15 84 100
2018/03/16 100 100
2018/03/17 100 100
2018/03/18 95 100
2018/03/19 93 100
2018/04/20 100 100
Notes from Meeting.
  • CMS will be migrating data from Castor to Echo with this expected to start in a few weeks time.